The value of a polynomial ax+by is 7 when x=3 and y=2, and is -3 when x=1 and y=2. Find a and b.
Answer by TimothyLamb(4379)

3a + 2b = 7
a + 2b = -3
a = -3 - 2b
---
3a + 2b = 7
3(-3 - 2b) + 2b = 7
-9 - 6b + 2b = 7
-16 = 4b
Answer:
b = -4
a = 5
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php